# Eye-tracked Virtual Reality: A Comprehensive Survey on Methods and Privacy Challenges

### v2

@@ IV-B3 Rendering Techniques
The quality of the 3D scene display is paramount for VR devices, as it directly impacts the overall immersive experience for users. Therefore, rendering with high resolution and high refresh rates in VR environments has been very important. However, several challenges exist regarding computational issues for real-time rendering, computational power, and realistic and natural depth perception. To overcome those, different rendering techniques utilizing gaze information have been proposed. In this section, we discuss rendering techniques in three categories based on their objectives: ensuring computational load and power efficiency, optimization of transmission efficiency, and scene quality enhancement. By exploring these categories, we aim to provide a deeper understanding of the current state of the art in rendering in VR and the potential for future developments. Table IV provides an overview of the papers in this section.
TABLE IV: A non-exhaustive overview of papers discussed in Section IV-B3.
@@
[204] Content aware foveated rendering using luminance contrast
@@
[206] Foveation, depth of field, and longitudinal chromatic aberration
@@
Saccade-based FR [209] Polynomial fitting for saccade landing position estimation
@@
[213] Gaze-based dynamic refocusing for gigapixel panoramas (GPP)
@@
[217] Low resolutions 2D scene processing and mapping to 3D scenes
@@
In this section, we first provide a comprehensive review of foveated rendering techniques. Subsequently, we discuss gaze-based methods with similar objectives. In a nutshell, foveated rendering approaches focus on graphical computations and rendering around gaze fixation in a high-quality manner and providing other parts of the scene in a relatively lower quality to reduce computation load without deteriorating the user experience. Figure 8 depicts illustrations of images with blurred peripheral regions.
To this end, Patney et al. [201] proposed a foveated rendering architecture to lessen computation power and reduce the number of shades in the scene. Their architecture generates a perceptual target image by adjusting the blur filter width according to retinal eccentricity to prevent temporal and spatial aliasing. The authors discovered that contrast enhancement could provide a larger tolerance for blurred region size through a user experience study. As a result, contrast enhancement was applied as a post-processing step to the image. The authors demonstrated that the proposed technique achieved a similar level of temporal stability compared to temporally filtered non-foveated images and validated it through frequency analyses, which showed alignment with the perceptual target.
Ye et al. [208] introduced rectangular mapping-based foveated rendering (RMFR), which renders scenes with non-uniform foveation levels based on eccentricity and scene complexity. RMFR offers superior visual quality compared to conventional foveated rendering methods while requiring minimal rendering cost. Meng et al. [203] proposed another foveated rendering approach, referred to as 3D-kernel foveated rendering (3D-KFR), which is combined with an eye-tracking-based acceleration algorithm. The proposed algorithm, explicitly designed for visualizing high-resolution microscopic 4D-Light fields with depth cues, provides rendering acceleration up to a factor of $7.28$ in HMDs.
Like the aforementioned works, Deng et al. [205] presented a gaze-contingent rendering approach with neural radiance fields (NeRF), which allows the rendering of 3D scenes with photo-realistic quality. However, this approach also requires heavy computation, which causes latency in the rendering process. To address this issue, the proposed technique involves encoding each retinal region, including the foveal, mid-periphery, and far-periphery regions, with varying levels of visual perception. According to the experiments, the authors indicated that the proposed method, incorporating eye-gaze information, reduces latency without compromising perceptual quality. Additionally, Meng et al. [202] took eye dominance into account and designed an approach providing more visual details to the dominant eye than the non-dominant eye to enhance the efficiency of foveated rendering. An image with a broader foveation area is rendered for the eye with higher ocular dominance. The authors conducted a study to estimate foveation level parameters individually for each eye, using the Miles test [228]. They showed a boost in the rendering performance of the devices with the proposed technique for the same level of perceived quality.
Most of the discussed foveated rendering algorithms consider the peripheral region and the sensitivity of human eyes to model foveated areas. In addition, Tursun et al. [204] introduced a content-aware foveated rendering method considering the luminance contrast of the exhibited content. This method can forecast the resolution parameters as a function of luminance patterns by processing low-resolution frames before performing high-resolution rendering. The authors’ user studies revealed that rendering performance increases while maintaining an imperceptible foveation layer. In a different study, Liu et al. [206] proposed a hybrid foveated rendering method incorporating foveation, field depth, and longitudinal chromatic aberration. This approach accounts for vergence and accommodation, visual acuity eccentricity, and color vision. The authors stated that the proposed model outperforms the state-of-the-art rendering techniques in frame rates while providing at least the same level of visual scene quality. In parallel, Kim et al. [207] presented a promising foveated rendering approach to deal with the high computational demand of ray tracing, which can be considered an advanced technique to achieve high visual quality and more realistic effects in scenes. The proposed foveated rendering approach includes a selective subsampling technique, which gradually decreases the rate in the peripheral region while preserving high sampling rates in the foveal region. This approach allows the practical use of ray tracing in HMDs by reducing the high power requirement.
A handful of the rendering techniques rely on fixation prediction, and such techniques may suffer from latency, which might result in the degradation of user experience because of the mismatch between actual gaze location and predicted location, especially during saccades. To this end, Arabadzhiyska et al. [209] provided an algorithm that forecasts the final position of saccades during the saccadic behavior and then renders the image considering the predicted landing location. To determine the final saccade location, the direction of the movements is predicted according to the previous gaze samples, while the saccade amplitude is predicted with polynomial fitting. The authors indicated that this technique could be integrated into the existing foveated rendering systems to improve performance. Similarly, Griffith and Komogortsev [210] proposed a data augmentation technique to improve the saccade landing point estimation with neural networks. Time-shifted imitations of the training data are used to boost the estimation reliability of the starting time of the saccadic movement. This augmentation improves the median accuracy of the predicted final saccade points for LSTMs and feed-forward neural networks.
Apart from the discussed foveated rendering techniques, qualitative or quantitative performance evaluation approaches are necessary to properly assess and compare the performance of the existing rendering techniques. In addition, the robustness and effectiveness of the introduced assessment methods could vary in different test scenarios. Hsu et al. [211] evaluated subjective assessment methods regarding efficiency and consistency for foveated rendering. Efficiency is assessed based on the required time to reach perceptual ratio convergence. Meanwhile, the distribution of individual quality of experience scores, which indicates user satisfaction, expectation, and perception levels, is used for consistency. The authors stated that no subjective evaluation method is superior to any other, but they still provide the research community with information on evaluation metrics so that researchers can decide on a metric that fits their specific needs.
@@
Figure 8: Foveated rendering provides a higher level of detail at the viewer’s gaze location while reducing the level of detail in the peripheral vision. We depict the gaze locations with blue circles in the figures.
@@
While we present a comprehensive discussion of various foveated rendering techniques designed to alleviate VR devices’ high power and GPU requirements and make them practically usable for real-world applications, several other techniques based on eye tracking address similar challenges in different ways. These are mainly based on gaze-based color manipulation and dimming techniques [212, 214]. To this end, Duinkharjav et al. [212] presented a real-time applicable power reduction method based on gaze information and color discrimination. The authors achieved a display power requirement reduction of up to 24% with minimal degradation in perceptual quality. Similarly, Wee et al. [214] proposed a system called FocusVR to overcome the high power requirements of VR and AR devices with advanced high-resolution displays. They enhanced the power consumption efficiency by combining techniques like intelligent dimming, vignettes, and color mapping. The eye-tracking capability of these devices is utilized to execute more aggressive dimming techniques with less impact on user experience. The authors reported that FocusVR could reduce display and system power consumption up to 80% and 50%, respectively. In addition to these works, Lyu et al. [213] proposed an approach using eye-tracking data to deal with high GPU requirements and lags in gigapixel panorama (GPP) displays in HMDs. A rendering technique supporting panning, tilting, zooming, and dynamic refocusing based on gaze information was proposed to display GPP scenes in HMDs. The introduced system can keep the frames per second (FPS) above 50 without a high-end GPU.
@@
Foveated rendering techniques can also be useful or customized to reduce the bandwidth requirements of the high-resolution 360-degree video streaming, especially for VR [215, 216]. In such setups, systems process eye-tracking data, and content providers transfer the data based on gaze position using a similar methodology employed in foveated rendering. Although different approaches exist, we depict the general framework for streaming solutions to optimize bandwidth efficiency in Figure 9.
@@
To this end, Li et al. [215] proposed a log-rectilinear transformation-based method using summed-area table filtering and off-the-shelf video codecs, that optimizes huge bandwidth requirements of the high-resolution live video streaming services for VR headsets. Log-rectilinear-based transformation provides about 10% efficiency in bandwidth usage compared to traditional log-polar transformation. It also reduced the flickering as measured by a metric suggested by Winkler et al. [229]. As an alternative, Lungaro et al. [216] presented a novel codec method, which intentionally causes some errors in the less-sensitive visual areas to support streaming video frames having regions with different qualities. The proposed solution offers an 83% improvement in the bandwidth requirement for high Quality of Service (QoS) levels over traditional solutions. Additionally, Chen et al. [217] proposed a gaze-contingent streaming approach that identifies the essential regions in 2D scenes and maps them to 3D scenes to improve rendering and streaming efficiency. The authors argued that the proposed system provides a better visual scene quality with temporal consistency than other systems.
@@
The aforementioned works, such as in foveated or streaming techniques, offer solutions to deal with or mitigate VR devices’ technical and computational limitations, such as issues regarding power consumption, computational load, and bandwidth requirements. In addition, achieving high scene quality and high levels of realism are other challenging concepts in VR, which are related to the visual acuity of virtual scenes. Most research in this field focuses on depth perception, which is related to the degree of realism. Therefore, in the following, we discuss the rendering techniques introduced to enhance realism, focusing mainly on the users’ depth perception. We also explore the proposed solutions for enhancing depth perception, which addresses the vergence-accommodation conflict (VAC).
@@
Mauderer et al. [218] presented a gaze-contingent technique allowing the production of more realistic 3D images with improved depth perception by modifying the blurring rate of the objects in a range of distances. The authors achieved higher levels of perceived depth and realism by employing focal points in combination with the gaze. Konrad et al. [219] proposed a rendering method based on ocular parallax exploiting small changes of images on the retina caused by eye rotation. The authors evaluated the ocular parallax’s visibility with experiments to measure perceived relative distance in the AR and VR scenes. They found that ocular parallax rendering combined with gaze contingency offers a more authentic experience and improved depth perception in VR.
@@
Another work accounting ocular parallax presented by Krajancih et al. [220] as a solution to the depth distortion in conventional rendering models. The introduced gaze-contingent stereo rendering method reduces the shape and disparity distortions and provides a consistent sense of depth while maintaining perceived relative distances between the rendered objects in a scene. Kellnhofer et al. [221] introduced another technique by using eye tracking as well as predicted gaze data to enhance depth perception while preserving the visual quality and user comfort. The authors’ model employs a controller executing unperceived local depth adjustments for the stereoscopic content to optimize the trade-off between depth reproduction and visual comfort, significantly enhancing the depth perception while preserving the visual quality.
@@
VAC, the mismatch between the vergence and accommodation points of the eye, is another challenge that can affect the perceived image quality, the sense of depth, and realism in near-eye displays. This conflict occurs due to changes in the shape of the eye lens during eye movements. In some VR displays, which are designed for a fixed distance, the degradation of perceived image quality can be attributed to the VAC. To address this challenge, Matsuda et al. [222] proposed the use of a phase-only spatial light modulator (SLM) made from the liquid crystal on silicon as a lens with a dynamically adaptive shape that is positioned between the display screen and the viewing optics. The authors presented a rendering technique that decomposes the scene and generates images for each display plane while considering the gaze location.
@@
Mercier et al. [223] proposed a computationally plausible decomposition method that utilizes gaze and head motions to align the layers of the multifocal displays in HMDs. The experiments, conducted on a multi-focal testbed designed for eye-tracking and accommodation measures, showed that eye tracking could be effectively utilized for plane alignment tasks. In addition, Konrad et al. [225] introduced an accommodation invariant display approach for near-eye displays, which makes scenes independent of the accommodation state of the eyes. This approach replaces traditional retinal blur with disparity and vergence to drive accommodation. Batmaz et al. [227] examined the effect of VAC on user performance during a hand-pointing task within arm’s reach using multi-focal and single-focal VR displays. The authors concluded that multi-focal displays are better than single-focal displays in terms of both time and error rate, corresponding to the percentage of missed targets during virtual hand interaction tasks. More recently, Aizenman et al. [224] systematically explored eye movement and binocular disparities when wearing an HTC Vive Pro Eye. The authors proposed an optimal screen distance to minimize the adverse effects of the VAC by analyzing the distribution of fixation distances. However, the accuracy of eye tracking can pose challenges for these solutions. To this end, Dunn [226] reported that the minimum eye-tracking accuracy required to implement varifocal displays in VR is $1.444^{\circ}$.
@@ IV-C Understanding Human Cognition, Visual Attention, and Perception
Computer vision- and machine learning-based methods to identify eye regions and detect gaze mostly work online for rendering and real-time user support. In addition, researchers often collect eye-tracking datasets from VR to analyze and understand human behaviors from a cognitive and visual attention perspective. To address those, in this section, we review the research on eye-tracked VR concerning understanding human behaviors. We first outline in Section IV-C1 how eye tracking can be helpful to understand human cognition in VR and during HMD use. Then, we explore the works investigating the relationship between human visual attention and gaze behaviors in Section IV-C2. Later, in Section IV-C3, we provide the research related to human perception of VR spaces, considering perceived realism and cybersickness.

### v3

@@ IV-B3 Rendering Techniques
The quality of the 3D scene display is paramount for VR devices, as it directly impacts users’ overall immersive experiences. Therefore, rendering with high resolution and high frame rates in VR has been very important. However, several challenges exist regarding computational issues for real-time rendering, computational power, and realistic and natural depth perception. To overcome those, different rendering techniques utilizing gaze information have been proposed. In this section, we discuss rendering techniques in three categories based on their objectives: ensuring computational load and power efficiency, optimizing transmission efficiency, and enhancing scene quality. By exploring these categories, we aim to provide a deeper understanding of the current state of the art in rendering in VR and the potential for future developments. Table IV provides an overview of the papers in this section.
TABLE IV: A broad overview of papers in Section IV-B3.
@@
[178] Content-aware foveated rendering
@@
[180] Foveated depth of field rendering
@@
Saccade-based FR [183] Polynomial fitting for saccade landing point estimation
@@
[187] Gaze-based dynamic refocusing for gigapixel panoramas
@@
[191] Low-resolution 2D scene processing and 3D mapping
@@
In this section, we provide a comprehensive review of foveated rendering techniques and discuss gaze-based methods with similar objectives. In a nutshell, foveated rendering approaches focus on high-quality graphical computations and rendering around gaze fixation, while providing other parts of the scene in a relatively lower quality to reduce computational load without deteriorating the user experience. Figure 7 depicts image illustrations with blurred peripheral regions.
@@
Figure 7: Foveated rendering provides a higher level of detail at the viewer’s gaze location while reducing the level of detail in the peripheral vision. We depict exemplary gaze locations with blue circles.
To this end, [175] proposed a foveated rendering architecture to lessen computational power and reduce the number of shades in the scene. This architecture generates a perceptual target image by adjusting the blur filter width according to retinal eccentricity to prevent temporal and spatial aliasing. The authors discovered that contrast enhancement could provide a larger tolerance for blurred region size through a user study. As a result, contrast enhancement was applied as a post-processing step to the image. The proposed technique achieved a similar level of temporal stability compared to temporally filtered non-foveated images. Another approach, rectangular mapping-based foveated rendering (RMFR), renders scenes with non-uniform foveation levels based on eccentricity and scene complexity [182]. RMFR offers superior visual quality compared to conventional foveated rendering methods while requiring minimal rendering cost. [177] proposed 3D-kernel foveated rendering (3D-KFR), which is combined with an eye-tracking-based acceleration algorithm. The proposed algorithm, designed for visualizing high-resolution microscopic 4D light fields with depth cues, provides rendering acceleration up to a factor of $7.28$ in HMDs.
Like the works above, [179] presented a gaze-contingent rendering approach with neural radiance fields (NeRF), which allows the rendering of 3D scenes with photo-realistic quality. However, this approach also requires heavy computation, which causes latency in the rendering process. To address this issue, the proposed technique involves encoding each retinal region, including the foveal, mid-periphery, and far-periphery regions, with varying levels of visual perception. Incorporating eye-gaze information reduces latency without compromising perceptual quality. Additionally, [176] took eye dominance into account and designed an approach that provides more visual details to the dominant eye than to the non-dominant eye, thereby enhancing the efficiency of foveated rendering. An image with a broader foveation area is rendered for the eye with higher ocular dominance, and it showed a boost in the rendering performance of the devices using the proposed technique at the same level of perceived quality.
Most discussed foveated rendering algorithms consider the peripheral region and the sensitivity of human eyes to model foveated areas. In addition, [178] introduced a content-aware foveated rendering method considering the luminance contrast of the exhibited content, which can forecast the resolution parameters as a function of luminance patterns by processing low-resolution frames before performing high-resolution rendering. Evaluations revealed that rendering performance increases while maintaining an imperceptible foveation layer. [180] proposed a hybrid foveated rendering method incorporating foveation, field depth, and longitudinal chromatic aberration. This approach accounts for vergence and accommodation, visual acuity eccentricity, and color vision, and it outperforms the state-of-the-art rendering techniques in frame rates while providing at least the same level of visual quality. In parallel, [181] presented a foveated rendering approach to address the high computational demands of ray tracing. The proposed approach incorporates a selective subsampling technique that gradually decreases the sampling rate in the peripheral region while maintaining high sampling rates in the foveal region. This approach allows the practical use of ray tracing in HMDs by reducing the high power requirement.
A handful of rendering techniques rely on fixation prediction, and such techniques may suffer from latency, which can result in degraded user experience due to a mismatch between the actual gaze location and the predicted location, especially during saccades. To this end, [183] provided an algorithm that forecasts the final position of saccades and then renders the image considering the predicted landing location. To determine the final saccade location, the direction of the movement is predicted based on the previous gaze samples, while the saccade amplitude is predicted using polynomial fitting, which can be integrated into existing foveated rendering systems to improve performance. Similarly, [184] proposed a data augmentation technique to enhance the saccade landing point estimation with neural networks. Time-shifted imitations of the training data are used to enhance the reliability of the estimation of the starting time of the saccadic movement. This augmentation improves the median accuracy of the predicted final saccade points for LSTMs and feed-forward neural networks.
Apart from the discussed foveated rendering techniques, qualitative or quantitative performance evaluation approaches are necessary to properly assess and compare the performance of the existing rendering techniques. Additionally, the robustness and effectiveness of the introduced assessment methods may vary across different test scenarios. [185] evaluated subjective assessment methods regarding efficiency and consistency for foveated rendering. Efficiency is assessed based on the required time to reach perceptual ratio convergence. Meanwhile, the distribution of individual quality of experience scores, which indicate user satisfaction, expectations, and perceptions, is used for consistency. The authors stated that no subjective evaluation method is superior to any other; however, they still provide the research community with information on evaluation metrics, allowing researchers to select a metric that best fits their specific needs.
@@
While we present various foveated rendering techniques designed to alleviate the high power and graphics processing unit (GPU) requirements of VR devices and make them practically usable for real-world applications, several other techniques based on eye tracking address similar challenges [186, 188]. To this end, [186] presented a real-time applicable power reduction method based on gaze and color discrimination, and achieved a display power requirement reduction of up to 24% with minimal degradation in perceptual quality. [188] proposed FocusVR to overcome the high power requirements of VR and AR devices with advanced high-resolution displays. They enhanced power consumption efficiency by combining techniques such as intelligent dimming, vignettes, and color mapping. The eye-tracking capability of these devices is utilized to execute more aggressive dimming techniques with less impact on user experience. FocusVR could reduce display and system power consumption up to 80% and 50%, respectively. In addition, [187] proposed an approach using eye tracking to address the high GPU requirements and lags associated with gigapixel panorama (GPP) displays in HMDs. A rendering technique supporting panning, tilting, zooming, and dynamic refocusing based on gaze was proposed to display GPP scenes in HMDs, and this technique can keep the frames per second (FPS) above 50 without a high-end GPU.
@@
Foveated rendering techniques can also be helpful or customized to reduce the bandwidth requirements of the high-resolution 360-degree video streaming, especially for VR [189, 190]. In such setups, systems process eye-tracking data, and content providers transfer the data based on gaze position analogous to foveated rendering. Despite different approaches, Figure 8 depicts the general framework for streaming solutions to optimize bandwidth efficiency.
@@
To this end, [189] proposed a log-rectilinear transformation-based method using summed-area table filtering and off-the-shelf video codecs, that optimizes huge bandwidth requirements of the high-resolution live video streaming services for VR HMDs. Log-rectilinear-based transformation achieves approximately 10% efficiency in bandwidth usage compared to the traditional log-polar transformation. Alternatively, [190] presented a novel codec method that intentionally introduces errors in less-sensitive visual areas to support streaming video frames with regions of varying quality. This method offers an 83% improvement in bandwidth requirement for high Quality of Service (QoS) levels compared to traditional solutions. Additionally, [191] proposed a gaze-contingent streaming approach that identifies the essential regions in 2D scenes and maps them to 3D scenes to enhance rendering and streaming efficiency. This approach is argued to provide better visual scene quality with temporal consistency compared to others.
@@
Foveated rendering or streaming techniques offer solutions to address or mitigate VR devices’ technical and computational limitations, such as power consumption, computational load, and bandwidth requirements. In addition, achieving high scene quality and high levels of realism are other challenging concepts in VR, which are related to the visual acuity of scenes. Most research in this direction focuses on depth perception, which is associated with the degree of realism. Therefore, in the following, we discuss the rendering techniques introduced to enhance realism, with a primary focus on users’ depth perception. We also explore the proposed solutions for improving depth perception, which addresses the vergence-accommodation conflict (VAC).
@@
[192] presented a gaze-contingent technique that allows for the production of realistic 3D images with improved depth perception by modifying the blurring rate of objects at varying distances. The authors achieved higher levels of perceived depth and realism by employing focal points in combination with the gaze. [193] proposed a rendering method based on ocular parallax exploiting small changes of images on the retina caused by eye rotation. The evaluations of ocular parallax’s visibility for measuring perceived relative distance in AR and VR scenes indicated that ocular parallax rendering combined with gaze contingency offers a more authentic experience and improved depth perception in VR.
@@
Another work on ocular parallax that focuses on gaze-contingent stereo rendering that reduces shape and disparity distortions, providing a consistent sense of depth while maintaining perceived relative distances between the rendered objects in a scene [194]. [195] introduced another technique that utilizes eye tracking and predicted gaze data to enhance depth perception while preserving visual quality and user comfort. The proposed method employs a controller that executes unperceived local depth adjustments for stereoscopic content to optimize the trade-off between depth reproduction and visual comfort, significantly enhancing depth perception while preserving visual quality.
@@
VAC, the mismatch between the vergence and accommodation points of the eye, is another challenge that can affect the perceived image quality, the sense of depth, and realism in near-eye displays. This conflict arises due to changes in the shape of the lens of the eye during eye movements. In some VR displays, which are designed for a fixed distance, the degradation of perceived image quality can be attributed to the VAC. To address this, [196] proposed the use of a phase-only spatial light modulator (SLM) made from the liquid crystal on silicon as a lens with a dynamically adaptive shape that is positioned between the display screen and the viewing optics. The authors presented a rendering technique that decomposes the scene and generates images for each display plane while considering the gaze location.
@@
Furthermore, [197] proposed a computationally plausible decomposition method that utilizes gaze and head motions to align the layers of the multifocal displays in HMDs. The evaluations indicated that eye tracking could be effectively utilized for plane alignment tasks. In addition, [199] introduced an accommodation invariant display approach for near-eye displays, which makes scenes independent of the accommodation state of the eyes. This approach replaces traditional retinal blur with disparity and vergence to drive accommodation. [201] examined the effect of VAC on user performance during a hand-pointing task within arm’s reach using multi-focal and single-focal VR displays, and concluded that multi-focal displays are better than single-focal displays in terms of both time and error rate, corresponding to the percentage of missed targets during virtual hand interaction tasks. More recently, [198] systematically explored eye movement and binocular disparities when wearing an HMD and proposed an optimal screen distance to minimize the adverse effects of VAC by analyzing fixation distance distributions. However, eye-tracking accuracy can pose challenges for such solutions. To this end, [200] reported that the minimum required eye-tracking accuracy for varifocal displays in VR is $1.444^{\circ}$.
@@ IV-C Understanding Human Cognition, Visual Attention, and Perception
Computer vision and machine learning methods for identifying eye regions and detecting gaze mostly work online for rendering or providing real-time user support. Researchers also often collect eye-tracking data from VR to analyze and understand human behaviors in an offline fashion. We first outline in Section IV-C1 how eye tracking can aid in understanding human cognition in VR and during HMD use. Then, we explore the works investigating the relationship between human visual attention and gaze behaviors in Section IV-C2. Later, in Section IV-C3, we provide the research on human perception of VR, considering perceived realism and cybersickness.